OpenStudy (anonymous):

what is the distance between the points (2,6) and (5,2)

OpenStudy (zehanz):

Use the distance formula, which is just the Pythagorean Theorem for points in the plane: If (a,b) and (c,d) are two points, then distance = sqrt((a-c)²+(b-d)²)

OpenStudy (anonymous):

thanks going to try it put out now

OpenStudy (zehanz):

Let me hear what you get ;)

OpenStudy (anonymous):

3

OpenStudy (zehanz):

I got: sqrt((2-5)²+(6-2)²)=sqrt(9+16)=sqrt(25)=5

OpenStudy (anonymous):

oh wow, for some reason i had the numbers line up different.

OpenStudy (zehanz):

sh*t happens ;)

OpenStudy (anonymous):

im really bad at math

OpenStudy (anonymous):

hey do you mind check my other math problems

OpenStudy (zehanz):

Tomorrow...I'm going to sleep now, It's midnight here!